OpenStudy (anonymous):

How do you simplify the expression -4 to the exponent of 2

OpenStudy (anonymous):

You square -4

OpenStudy (anonymous):

re-write it twice\[(a)^2=a*a\]

OpenStudy (anonymous):

(-4)^2=(-4)*(-4)=(-1)*(-1)*16=16

OpenStudy (anonymous):

Multiply -4 by -4